Evaluate the integral of sin2x dx/(2+sin² x)


1
Expert's answer
2021-03-29T14:01:54-0400

Ans:-

"\int\dfrac{Sin2x}{2+sin^2x}dx"


"Sin2x=2\times(Sinx)\times(Cosx)"

put "2+Sin^2x=t \Rightarrow 2Sinxcosxdx=dt"


"\int\dfrac{dt}{t}=lnt +C"


"\Rightarrow ln(2+Sin^2x)+C"